the mass of an electron is 9.11x10^-31 kg. charge on the electrcon is 1.69x10^-19 C and g = 9.8 m/s^2 then the intensity of eletric field required to balance weight of an electron is?
1
Expert's answer
2020-08-02T15:07:26-0400

The weight of electron is:


"P = mg"

where "m = 9.11\times 10^{-31}kg" and "g = 9.8 m\/s^2".

The electric force acting on electron in the field "E" is the following:


"F = qE"

where "q = 1.69\times 10^{-19}C". The balance condition:


"P = F\\
mg = qE\\"

Expressing the intensity of eletric field, obtain:


"E = \dfrac{mg}{q} = \dfrac{9.11\times 10^{-31}\cdot 9.8}{1.69\times 10^{-19}} \approx 5.28 \times10^{-11} N/q"

Answer. 5.28x10^-11 N/q.